W start up Ok.

Secondary Start up, Ok, with the time to power on and setflatAO.

Testing all the system with a close loop.
Fault during the close loop, the system gives ok and all the status AOSet Execution Success, BUT the CCD39 viewer goes to NOT LIVE and looking the Master Diagnostic we have the following error message: "MASTDIAGN01 |WAR| 2032|2009-10-13 10:28:30.935404| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 2033|2009-10-13 10:28:31.255480| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 2034|2009-10-13 10:28:31.575567| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 2035|2009-10-13 10:28:31.895669| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 2036|2009-10-13 10:28:31.950638| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 2037|2009-10-13 10:28:32.270685| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 2038|2009-10-13 10:28:32.590752| OPTICAL-LOOP > Wrong filotto. Frame skipped."

Reseting the BCU39 using the option BCU39 Reseting didn't solve the problem. Looking the TX fiber, seems that is not trasmitting signal ( no light).

Power off and Power cycle OFF the WFS UNIT.

Power on the Master is giving the same error message: "MASTDIAGN01 |WAR| 4947|2009-10-13 10:41:59.384184| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 4948|2009-10-13 10:41:59.704269| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 4949|2009-10-13 10:42:00.024330| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 4950|2009-10-13 10:42:00.076838| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 4951|2009-10-13 10:42:00.396876| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 4952|2009-10-13 10:42:00.716967| OPTICAL-LOOP > Wrong filotto. Frame skipped.
MASTDIAGN01 |WAR| 4953|2009-10-13 10:42:01.037022| OPTICAL-LOOP > Wrong filotto. Frame skipped.
..........."

Power Off- On and restart the software. This operation doens't show the error message in the Master Diagnostic. Looking the log we have the right message:
"NF| 67|2009-10-13 10:44:19.531351| MAIN > Created RTDB var R.MASTDIAGN01.CLOSELOOP.CUR
MASTDIAGN01 |INF| 68|2009-10-13 10:44:19.531393| MAIN > Open/Close optical-loop RTDB vars support installed...
MASTDIAGN01 |INF| 69|2009-10-13 10:44:19.531413| MAIN >
MASTDIAGN01 |INF| 70|2009-10-13 10:44:19.531424| MAIN > ------------------- STARTING MASTER DIAGNOSTIC SYSTEM --------------------
MASTDIAGN01 |ALW| 71|2009-10-13 10:44:19.531447| MAIN > Real-time priority (99) succesfully set !
MASTDIAGN01 |INF| 72|2009-10-13 10:44:19.531457| MAIN > > STARTING OPTICAL LOOP DIAGNOSTIC...
MASTDIAGN01 |ALW| 73|2009-10-13 10:44:19.531547| MAIN > Real-time priority (99) succesfully set !
MASTDIAGN01 |ALW| 74|2009-10-13 10:44:19.531560| BCU39 > FramesDownloader BCU39: READY TO DOWNLOAD FRAMES OF SIZE 19248 BYTES (0 = disabled)...
MASTDIAGN01 |ALW| 75|2009-10-13 10:44:19.531585| MAIN > Real-time priority (99) succesfully set !
MASTDIAGN01 |ALW| 76|2009-10-13 10:44:19.531597| ADSEC_STREAM > FramesDownloader ADSEC_STREAM: READY TO DOWNLOAD FRAMES OF SIZE 8120MASTDIAGN01 |ALW| 77|2009-10-13 10:46:43.478008| OPTICAL-LOOP > Frequency: 0.69 Hz (100)
MASTDIAGN01 |ALW| 78|2009-10-13 10:46:45.000403| OPTICAL-LOOP > Frequency: 65.69 Hz (200)
MASTDIAGN01 |ALW| 79|2009-10-13 10:46:46.522760| OPTICAL-LOOP > Frequency: 65.69 Hz (300)
........
"
AdSEc Controlm Gui: Shows all the options less than the Gain vector and Recover Fail not availables . AdsecArbitrator StatusAORunning. Last command: Set Gain Execution success.
Stop and Start the Arbitratorchange the status to AORunning, BUT ALL THE BUTTONS ARE NOT AVAILABLE
THE BCU39 STOP TO TRANSMIT SLOPES, LOOSING THE COMMUNICATION. CHECKING THE MIRROR CONTROLLER AND THERE IS A ERROR DURING THIS TIME. THE BCU RESTARTED BY ITS OWN, AND IT WAS POSSIBLE TO OPEN AND CLOSE THE LOOP AGAIN. CHANGE THE NUMBER OF PACKETS WHICH IS CHECKING FROM 2 TO 5.
RIP THE SECONDARY WITH A OPEN LOOP ACTION.

.................CLOSE LOOP BIN 3 6L/D 200Hz. low flux. Modes:66

CLOOP <_125110>
_125248
IRTC dark = <20091013_140150>

IR exp =300000 us
IR n frames = 100
Bin = <3>
LoopHz = <200>
Pup = <tracking#>
Dist =<ovfreq800 / oversampling>
Gain = <0.5>
Mod = 6 l/D
Rec = <20090930_120512>
SN = zero/....
FW1 = <600>
FW2 =
Phot = 40 phot/subap
Staged: <On>
Nota = <................>

CLOOP <_125443>
_125526
IRTC dark = <20091013_140150>

IR exp =300000 us
IR n frames = 100
Bin = <3>
LoopHz = <200>
Pup = <tracking#>
Dist =<ovfreq800 / oversampling>
Gain = <0.5>
Mod = 6 l/D
Rec = <20090930_120512>
SN = zero/....
FW1 = <600>
FW2 =
Phot = 40 phot/subap
Staged: <OFF>
Nota = <................>

CLOOP <_125823>
_125904
IRTC dark = <20091013_140150>

IR exp =300000 us
IR n frames = 100
Bin = <3>
LoopHz = <200>
Pup = <tracking#>
Dist =<ovfreq800 / oversampling>
Gain = <0.9>
Mod = 6 l/D
Rec = <20090930_120512>
SN = zero/....
FW1 = <600>
FW2 =
Phot = 40 phot/subap
Staged: <OFF>
Nota = <................>

IRTC_DARK 500000us 20091013_140602

---+++ IMPLEMENTATO OSS per l'esapodo -- AlfioPuglisi - 2009-10-13 La GUI dell'OSS sembra funzionare

Misure CLOOP BIN3 mag 14.4
173754 ........ 174419

-- AlfioPuglisi - 2009-10-13
Topic revision: r5 - 14 Oct 2009, GuidoAgapito
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback